OpenStudy (anonymous):

6 over the 5th root of 16

OpenStudy (anonymous):

\[\frac{ 6 }{ \sqrt[5]{16} }\]

OpenStudy (anonymous):

i had the question typed all wrong

OpenStudy (anonymous):

Similar to the last question you need to rationalize that denominator! The difference being that in order to rationalize it you will need to multiply a few more times. So this time we will multiply the fifth root of 16 4 times.

OpenStudy (anonymous):

i simplified the 16 to \[2^{4}\]

OpenStudy (anonymous):

\[\frac{ 6 }{ \sqrt[5]{16} } * \frac{ \sqrt[5]{16} }{ \sqrt[5]{16} }* \frac{ \sqrt[5]{16} }{ \sqrt[5]{16} }* \frac{ \sqrt[5]{16} }{ \sqrt[5]{16} }* \frac{ \sqrt[5]{16} }{ \sqrt[5]{16} }\]

OpenStudy (anonymous):

\[\frac{ 6 * \sqrt[5]{65536} }{ 16 }\]

OpenStudy (anonymous):

You can simplify 16 to 2^4 but that's not quite simplest form, we usually try to get rid of exponents because they muddy things :)

OpenStudy (anonymous):

So from here how would you simplify further?